# A Comprehensive Survey with Critical Analysis for Deepfake Speech Detection

### v1

@@ IV-A Offline data augmentation
Analysis: Data augmentation involves generating variations of the original data to increase the size of DSD datasets, which enhances the robustness and generalization capabilities of machine learning models. Since this step is applied to original audio utterances before the training process, it can be referred to as offline data augmentation. As shown in Tables V, VI, and VII, offline data augmentation methods can be separated into two main groups, referred to as compression and distortion. The compression methods involve compress and decompress algorithms, mainly using audio codec techniques. A codec, short for ‘coder-decoder’, is software used to compress and decompress digital audio. Among these methods, MP3, AAC, OGG, G.7XX, and Opus formats are commonly applied. Codec data augmentation helps simulate these real-world conditions through various compression schemes (e.g., phone calls, music streaming, or online video playback on applications such as Facebook, WhatsApp, etc.). Since different codecs use various compression and decompression algorithms, they impact audio-related factors such as signal-to-noise ratio (SNR), high-frequency formants, energy loss, sample rate, bit depth, and bitrate in distinct ways. This suggests that if there are subtle differences between real and fake speech in these aspects, generating diverse audio utterances using different codecs can be an effective approach for distinguishing between them.
@@
Discussion: Although compression methods and distortion methods present different approaches to generate more audio data, none of the papers has compared, analyzed, and indicated if one of the approaches is superior in the DSD task. Indeed, the statistical information in Fig 6 indicates that the state-of-the-art DSD systems using offline distortion augmentation and offline compression augmentation are equal.
@@
Our contribution: Given the analysis and the existing concerns above, our work mainly contributes:
  • To evaluate the role and the effect of the online and offline data augmentation methods, we conducted extensive experiments in this paper. Based on our findings, we identify data augmentation techniques that are compatible with DSD systems. In particular, we compare the performance of codec-based methods with the Mixup [189] and SpecAugment [190] in this paper.
@@ IV-B Feature extraction
Regarding feature extraction using audio embeddings from pre-trained models, although these pre-trained models are effective for many audio tasks, using them for deepfake detection presents several challenges. Firstly, as pre-trained models are initially trained for upstream tasks such as speech-to-text, speaker identification, emotion detection, etc, that focus on different aspects (i.e., speech-to-text or emotion detection), the audio melody and harmony (i.e. emotion detection), or distinct frequencies (i.e., speaker identification), embeddings can fail to capture subtle artifacts specific to synthesized speech. Secondly, audio deepfakes are generated to closely mimic real speech, they often have the same formants, pitch, and rhythm as real audio, especially when generated by advanced deep-learning-based speech generation systems. Additionally, the use of pre-trained models can add complexity due to their large network architectures.
@@
Our contribution: Given the analysis and the discussion above, our work mainly contributes:
@@ IV-C Classification models
Analysis: Early models proposed for DSD task approached conventional machine learning algorithms. For example, 9 over 16 submitted systems in ASVspoofing 2015 challenge [191] extract MFCC feature (i.e. Systems A, B, E, G, H, I, N, O, and P in [191]). Then, various machine learning-based models such as Mahalanobis distance measurement, Gaussian-based model (GMM), support vector machine-based models (SVM, SVM-RBF), or fusion models (GMM and SVM) are used to explore MFCC features. However, recent DSD systems as shown in Table VVIVII present a wide range of neural network architectures due to the powerful deep learning techniques. Recently proposed deep neural networks for the DSD task can be separated into four main approaches. The first approach, which focuses on exploring spatial features, leverages convolutional-based network architectures (CNN). Among the CNN-based networks, Resnet, LCNN, and RawNet architectures are widely approached. ResNet and LCNN are used to explore spectrogram-based features such as LFCC [127], CQT [132], and MEL [145]. Meanwhile, RawNet architectures are normally combined with SincNet layer [193] to learn raw audio [126, 127, 129, 130, 133, 145]. The second approach, which focuses on exploring the temporal features, presents recurrent neural network (RNN) based architectures. For example, LSTM-based networks, TDNN, or ECAPA-TDNN are proposed in [130, 126][158, 129] and [168], respectively. As shown in Table VVIVII, RNN-based networks have not been popularly applied for the DSD task compared to the CNN-based architectures. The third approach involves combining both convolutional layers and recurrent layers to explore both temporal and spatial features, referred to as hybrid network architectures. In particular, recurrent network-based layers such as LSTM, GRU are combined with CNN-based layers to perform convolutional-recurrent neural network (CRNN) architectures [126, 168, 165].
@@
Our contribution: Given the analysis and the discussion above, our work mainly contributes:
@@
  • Given extensive experiments on different input features and various network architectures, we propose an ensemble model that is competitive to the state-of-the-art DSD systems
@@ IV-D Loss function and training strategy
From Table VVIVII, it can be seen that most proposed models use a single loss function. Statistics of the individual loss functions are also presented in Fig. 7. As shown in Fig. 7, the cross entropy (CE) based losses (e.g. Binary Cross Entropy (BCE), Weight Cross Entropy (WCE), etc.) and Softmax-based losses (e.g., Additive-Margin-Based Softmax (AM-Softmax), Angular-Margin-Based Softmax (A-Softmax), etc.) present the most popular loss functions. Some models combine different loss functions. For example, CE and Contrastive loss were used in [169]. Similarly, authors in [165] combined three loss functions of Cross Entropy, Triplet loss, and Adversarial loss. Some papers such as [159] and [160] compared the DSD performance between large margin cosine loss (LMC loss), and A-Softmax loss functions or between OC-Softmax, MSE for P2SGrad loss functions, respectively.
@@ V-A Our motivation
  • We conduct extensive experiments to evaluate different inputs and various network architectures. Given the comparison, we indicate which input features, network architectures, combination of input features, and network architectures have the potential to be further explored. We then propose the best DSD ensemble system which is competitive to the state-of-the-art systems.
@@ V-B Selected datasets and evaluating metrics
As the trade-off among the number of utterances, the deep-learning-based fake speech generation systems, the original/real human speech resource shown in Fig. 2 and the comprehensive analysis in Section III, we decide to use ASVspoof 2019 (LA Task) to evaluate the effect of data augmentations, different types of input features, and various network architectures.
We obey the ASVspoof 2019 (LA Task) challenge, then use the Equal Error Rate (ERR) as the main metric for evaluating proposed models. We also report the Accuracy, F1 score, and AUC score to compare the performance among evaluating models.
@@ V-C Proposed systems and experimental settings
Back-end classification models: This paper proposes three main approaches for back-end classification models: the end-to-end deep learning approach, the transfer learning approach, and the audio-embedding deep learning approach. Regarding the end-to-end deep learning approach, four models of CNN-based model, SinC-CNN model (CNN architecture in SinC-CNN model reused from CNN-based model), RNN-based model, and C-RNN-based model are evaluated with the detailed configuration in Table VIII. The Sinc-CNN model proves powerful for raw audio input and has been widely used as the survey in Section IV Meanwhile, CNN-based models are commonly used and effectively capture and learn spectral features. We also use RNNs to focus on detecting natural sequential patterns that can be disrupted in synthetic audio [200] (e.g., temporal coherence, prosodic features such as rhythm, stress, and intonation). Consequently, based on the idea of combining both spectral features and temporal features, we use C-RNN-based model to distinguish characteristics of real and fake audio utterances.
@@
For the audio-embedding deep learning approach, the state-of-the-art audio pre-trained models of Whisper [150], Seamless [202], Speechbrain [203], and Pyanote [204, 205] are leveraged.
@@ V-D Experimental results and discussion
Evaluation of data augmentation methods: Considering the performance of online and offline data augmentation methods as shown in systems A1 (no data augmentation), A2 (online data augmentation with codec), A3 (offline data augmentation with Mixup and SpecAugment), and A4 (both online and offline data augmentation), it can be seen that the offline data augmentations of Mixup and SpecAugment are appropriate for DSD task. Notably, the combination of online and offline data augmentations does not help to enhance the DSD task performance compared with only using online data augmentation.
@@
Evaluation of input features: Considering the efficacy of raw audio and six types of spectrograms in systems from B1 to B7, STFT outperforms the raw audio and other spectrograms. Models B2, B5, and B7 achieve the best ERR score of 0.08 while the combination of STFT & LF obtains slightly better accuracy and F1 score of 0.88 and 0.9, respectively. This indicates that STFT and applying filters such as Linear Filter or Gammatone filter are suitable for isolating specific frequency bands in classification algorithms.
@@
Multiple deep learning approaches: Regarding the end-to-end deep learning approach from A1 to C2, CNN systems outperform RNN or C-RNN systems. Indeed, using the same input feature of STFT+LFCC, RNN and C-RNN approaches (C1 and C2 systems) obtain ERR scores of 0.14 and 0.17, which is significantly worse than CNN system (A3 or B2 or B7) with the best score of 0.08. This indicates that the specific patterns indicative of deepfake audio might not be primarily temporal but rather frequency in the spectrogram representation. Regarding the finetuning approach (D1 to D10), Convnext-Tiny stands out as the best system with competitive EER scores of 0.075. Meanwhile, the embedding-based approach (E1 to E4) achieves the best EER scores of 0.10 from the pre-trained Whisper model. This suggests the potential of these approaches when choosing the appropriate networks for enhancement.
@@
Ensembles: Given the performance of individual input features and network architecture, we conduct extensive experiments to evaluate a wide range of ensemble models. First, ensembles of STFT, CQT, and WT spectrograms are evaluated, indicating the best EER score of 0.06 from the combination of STFT and CQT (B2+B3). Then, ensembles of spectrogram with different filter banks (MEL, LF, GAM) are also evaluated, resulting in the best score of 0.065 from STFT+LF and STFT+GAM (B5+B7). As a result, when an ensemble of CQT, STFT+LF, and STFT+GAM is conducted (B3+B5+B7), we can achieve the EER score of 0.05. Regarding the ensemble of network architecture, CNN and ConvNeXt-Tiny (B5+D7) help obtain the EER score of 0.07. Meanwhile, the combination of Whisper+MLP and ConvNeXt-Tiny (E1+D7) achieves the best EER score of 0.03. These results obtained from ensemble models lead to some conclusions:
  • Ensembling different input features or network architectures can significantly help to enhance the DSD task performance.
@@
  • Not all network architectures are appropriate for the DSD task. As the good performance obtained from ConvNeXt-Tiny, the depthwise convolution layer in this network architecture needs to be further investigated.
@@
  • Leveraging pre-trained models such as Whisper shows effectiveness. This again proves why the recent Encoder-Decoder network architectures using pre-trained models for the Encoder have been proposed popularly.

### v4

@@ IV-A Offline data augmentation
Analysis: Data augmentation involves generating variations of the original data to increase the size of DSD datasets, which enhances the robustness and generalization capabilities of machine learning models. Since this step is applied to original audio utterances before the training process, it can be referred to as offline data augmentation. As shown in Tables V, VI, and VII, offline data augmentation methods can be separated into two main groups, referred to as compression and distortion. The compression methods involve compress and decompress algorithms, mainly using audio codec techniques. A codec, short for ‘coder-decoder’, is a software used to compress and decompress digital audio. Among these methods, MP3, AAC, OGG, G.7XX, and Opus formats are commonly applied. Codec data augmentation helps simulate these real-world conditions through various compression schemes (e.g., phone calls, music streaming, or online video playback on applications such as Facebook, WhatsApp, etc.). Since different codecs use various compression and decompression algorithms, they impact audio-related factors such as signal-to-noise ratio (SNR), high-frequency formants, energy loss, sample rate, bit depth, and bitrate in distinct ways. This suggests that if there are subtle differences between real and fake speech in these aspects, generating diverse audio utterances using different codecs can be an effective approach for distinguishing between them.
@@
Discussion: Although compression methods and distortion methods present different approaches to generate more audio data, none of the papers has compared, analyzed, and indicated if one of the approaches is superior in the DSD task. Indeed, the statistical information in Fig 6 indicates that the number of state-of-the-art DSD systems using offline distortion augmentation and offline compression augmentation are equal.
@@
Our contribution: Given the analysis and the existing concerns above, we make the following main contributions:
  • To evaluate the role and the effect of the online and offline data augmentation methods, we conducted extensive experiments in this paper. Based on our findings, we identify data augmentation techniques that are compatible with DSD systems. In particular, we compare the performance of codec-based methods with the Mixup [189] and SpecAugment [190].
@@ IV-B Feature extraction
Regarding feature extraction using audio embeddings from pre-trained models, although these pre-trained models are effective for many audio tasks, using them for deepfake detection presents several challenges. Firstly, as pre-trained models are initially trained for upstream tasks such as speech-to-text, speaker identification, emotion detection, etc, that focus on different aspects (i.e., speech-to-text or emotion detection), the audio melody and harmony (i.e., emotion detection), or distinct frequencies (i.e., speaker identification), embeddings can fail to capture subtle artifacts specific to synthesized speech. Secondly, audio deepfakes are generated to closely mimic real speech, they often have the same formants, pitch, and rhythm as real audio, especially when generated by advanced deep-learning-based speech generation systems. Additionally, the use of pre-trained models can add complexity due to their large network architectures.
@@
Our contribution: Given the analysis and the discussion above, we make the following main contributions:
@@ IV-C Classification models
Analysis: Early models proposed for DSD task approached conventional machine learning algorithms. For example, 9 over 16 submitted systems in ASVspoofing 2015 challenge [191] extract MFCC feature (i.e. Systems A, B, E, G, H, I, N, O, and P in [191]). Then, various machine learning-based models such as Mahalanobis distance measurement, Gaussian-based model (GMM), Support vector machine-based models (SVM, SVM-RBF), or fusion models (GMM and SVM) are used to explore MFCC features. However, recent DSD systems as shown in Table VVIVII present a wide range of neural network architectures due to the powerful deep learning techniques. Recently proposed deep neural networks for the DSD task can be separated into four main approaches. The first approach, which focuses on exploring spatial features, leverages convolutional-based network architectures (CNN). Among the CNN-based networks, Resnet, LCNN, and RawNet architectures are widely used. ResNet and LCNN are used to explore spectrogram-based features such as LFCC [127], CQT [132], and MEL [145]. Meanwhile, RawNet architectures are normally combined with SincNet layer [193] to learn raw audio [126, 127, 129, 130, 133, 145]. The second approach, which focuses on exploring the temporal features, presents recurrent neural network (RNN) based architectures. For example, LSTM-based networks, TDNN, or ECAPA-TDNN are proposed in [130, 126][158, 129] and [168], respectively. As shown in Table VVIVII, RNN-based networks have not been popularly applied for the DSD task compared to the CNN-based architectures. The third approach involves combining both convolutional layers and recurrent layers to explore both temporal and spatial features, referred to as hybrid network architectures. In particular, recurrent network-based layers such as LSTM, GRU are combined with CNN-based layers to perform convolutional-recurrent neural network (CRNN) architectures [126, 168, 165].
@@
Our contribution: Given the analysis and the discussion above, we make the following main contributions:
@@
  • Given extensive experiments on different input features and various network architectures, we propose an ensemble model that is competitive to the state-of-the-art DSD systems.
@@ IV-D Loss function and training strategy
From Table VVIVII, it can be seen that most proposed models use a single loss function. Statistics of the individual loss functions are also presented in Fig. 7. As shown in Fig. 7, the cross entropy (CE) based losses (e.g., Binary Cross Entropy (BCE), Weight Cross Entropy (WCE), etc.) and Softmax-based losses (e.g., Additive-Margin-Based Softmax (AM-Softmax), Angular-Margin-Based Softmax (A-Softmax), etc.) present the most popular loss functions. Some models combine different loss functions. For example, CE and Contrastive loss were used in [169]. Similarly, authors in [165] combined three loss functions of Cross Entropy, Triplet loss, and Adversarial loss. Some papers such as [159] and [160] compared the DSD performance between large margin cosine loss (LMC loss), and A-Softmax loss functions or between OC-Softmax, MSE for P2SGrad loss functions, respectively.
@@ V-A Our motivation
  • We conduct extensive experiments to evaluate different inputs and network architectures. Given the comparison, we indicate which input features, network architectures, combination of input features, and network architectures have the potential to be further explored. We then propose the best DSD ensemble system that is competitive to the state-of-the-art systems.
  • To deeply analyze the role of data augmentation methods, input features, and network architectures, we evaluated proposed DSD systems within cross-dataset and cross-language settings.
@@ V-B Selected datasets and evaluating metrics
As the trade-off among the number of utterances, the deep-learning-based fake speech generation systems, the original/real human speech resource as shown in Fig. 2 and the comprehensive analysis in Section III, we decide to use ASVspoof 2019 (LA Task) to evaluate the effect of data augmentations, different types of input features, and various network architectures. Given the results on ASVspoof 2019 (LA Task), we obtain the best DSD systems which are then evaluated with ASVspoof 2021 (LA & DF Tasks) datasets for cross-dataset evaluation and with MLAAD dataset for cross-language evaluation.
We obey the ASVspoof 2019 (LA Task) and ASVspoof 2021 (LA & DF Tasks) challenges, then use the Equal Error Rate (ERR) as the main metric for evaluating proposed models. We also report the Accuracy, F1 score, and AUC score to compare the performance among evaluating models.
@@ V-C Proposed systems and experimental settings
Back-end classification models: This paper proposes three main approaches for back-end classification models: the end-to-end deep learning approach, the transfer learning approach, and the audio-embedding deep learning approach. Regarding the end-to-end deep learning approach, four models of CNN-based model, SinC-CNN model (e.g., SinC-CNN architecture is a combination of SinC layer and CNN architecture. The CNN architecture component is reused from CNN-based model), CNN-based model, RNN-based model, and C-RNN-based model are evaluated with the detailed configuration in Table VIII. The Sinc-CNN model proves powerful for raw audio input and has been widely used as the survey in Section IV Meanwhile, CNN-based models are commonly used and effectively capture and learn spectral features. We also use RNNs to focus on detecting natural sequential patterns that can be disrupted in synthetic audio [200] (e.g., temporal coherence, prosodic features such as rhythm, stress, and intonation). Consequently, based on the idea of combining both spectral features and temporal features, we use C-RNN-based model to distinguish characteristics of real and fake audio utterances.
@@
For the audio-embedding deep learning approach, the state-of-the-art audio pre-trained models of Whisper [150], Seamless [202], Speechbrain [203], and Pyannote [204, 205] are leveraged.
@@ V-D Experimental results and discussion
We first use ASVspoof 2019 (LA Task) to evaluate and indicate the best DSD systems. The comprehensive result comparison is described in Table X.
Evaluation of data augmentation methods on ASVspoof 2019 (LA Task): Considering the performance of online and offline data augmentation methods as shown in systems A1 (no data augmentation), A2 (online data augmentation with codec), A3 (offline data augmentation with Mixup and SpecAugment), and A4 (both online and offline data augmentation), it can be seen that the offline data augmentations of Mixup and SpecAugment are appropriate for DSD task on ASVspoof 2019 (LA Task) dataset. Notably, the combination of online and offline data augmentations does not help to enhance the DSD task performance compared with only using offline data augmentation.
@@
Evaluation of input features on ASVspoof 2019 (LA Task): Considering the efficacy of raw audio and six types of spectrograms in systems from B1 to B7, STFT outperforms the raw audio and other spectrograms. Models B2, B5, and B7 achieve the best ERR score of 0.08 while the combination of STFT & LF obtains slightly better accuracy and F1 scores of 0.88 and 0.9, respectively. This indicates that STFT and applying filters such as Linear Filter or Gammatone filter are suitable for isolating specific frequency bands in classification algorithms.
@@
Evaluate multiple deep learning approaches on ASVspoof 2019 (LA Task): Regarding the end-to-end deep learning approach from A1 to C2, CNN systems outperform RNN or C-RNN systems. Indeed, using the same input feature of STFT+LFCC, RNN and C-RNN approaches (C1 and C2 systems) obtain ERR scores of 0.14 and 0.17, which is significantly worse than CNN system (A3 or B2 or B7), with the best score of 0.08. This indicates that the specific patterns indicative of deepfake audio might not be primarily temporal but rather frequency in the spectrogram representation. Regarding the finetuning approach (D1 to D10), Convnext-Tiny stands out as the best system with competitive EER scores of 0.075. Meanwhile, the embedding-based approach (E1 to E4) achieves the best EER scores of 0.10 using the pre-trained Whisper model. This suggests the potential of these approaches when choosing the appropriate networks for further optimization.
@@
Evaluate ensemble methods on ASVspoof 2019 (LA Task): Given the performance of individual input features and network architecture, we conduct extensive experiments to evaluate a wide range of ensemble models. First, ensembles of STFT, CQT, and WT spectrograms are evaluated, indicating the best EER score of 0.06 from the combination of STFT and CQT (B2+B3). Then, ensembles of spectrogram with different filter banks (MEL, LF, GAM) are also evaluated, resulting in the best score of 0.065 from STFT+LF and STFT+GAM (B5+B7). As a result, when an ensemble of CQT, STFT+LF, and STFT+GAM is conducted (B3+B5+B7), we can achieve the EER score of 0.05. Regarding the ensemble of network architecture, CNN and ConvNeXt-Tiny (B5+D7) help obtain the EER score of 0.07. Meanwhile, the combination of Whisper+MLP, ConvNeXt-Tiny (E1+D7) or Whisper+MLP, CNN (E1+B5) achieves the best EER score of 0.03.
We continue evaluating cross-datasets on ASVspoof 2021 (LA & DF Tasks) [23] and cross-languages on MLAAD dataset [18]. For the cross-dataset evaluation, the evaluation sets of ASVspoof 2021 (LA & DF Tasks) [23] are tested with the DSD models which were trained and evaluated on ASVspoof 2019 (LA Task) in advance from Table X. Regarding cross-language evaluation, we only select pairs of utterances from four languages (e.g. French, Spanish, Italian, and German). A pair of utterances presents the original utterance and a deepfake utterance with the same transcription. Similar to the cross-dataset evaluation, pre-trained DSD systems on ASVspoof 2019 (LA Task) from Table X are used to verify the cross-language evaluation.
Data augmentation methods for cross-dataset evaluation on ASVspoof 2021 (LA & DF Tasks): As experimental results on the B5 system are shown in Table XI, it indicates that using the offline data augmentation of codec helps improve the DSD system performance on both ASVspoof 2021 LA and DF tasks. Significantly, codec helps enhance by 0.11 in terms of EER score in the ASVspoof 2021 LA task. The results also indicate that a combination of offline data augmenation (e.g., codec) and online data augmentation (e.g., Mixup and SpecAugment) are necessary to achieve a general DSD model to deal with the domain shift issue in cross-data evaluation.
Input features for cross-dataset evaluation on ASVspoof 2021 (LA & DF Tasks): Regarding the input features, three types of spectrograms (e.g., CQT, STFT+GAM, STFT+LF) which present the high performance on ASVspoof 2019 dataset are evaluated. In particular, STFT+LF (B5 system) outperforms CQT (B3 system) and STFT+GAM (B7 system). This indicates that a combination of STFT and linear filter is suitable for DSD task.
Network architecture for the cross-dataset evaluation on ASVspoof 2021 (LA & DF Tasks): The experimental results from B5 (STFT+LF, CNN), D7 (STFT+LF, ConvNeXt-Tiny) and E1 (Raw Audio, Whisper+MLP) systems indicate that leveraging pre-trained model (E1) significantly outperforms the others. This again proves and explains why more Encoder-Decoder architectures have been recently proposed for the DSD task (i.e., Encoder architectures leveraging pre-trained models such as Whisper or Wave2vec2.0). Regarding the ensemble methods, the combination of D7 and E1, which present CNN model trained from scratch and pre-trained Whisper model, achieves the best performance on both ASVspoof 2019 (LA Task) and ASVspoof 2021 (LA & DF Tasks). This also proves that the ensemble of network architectures is more effective than the ensemble of input features.
The results obtained from the evaluation on ASVspoof 2019 (LA Task) and ASVspoof 2021 (LA & DF Tasks) lead to some conclusions:
  • The results indicate a combination of offline data augmentation (codec) and online data augmentation (Mixup, SpecAugment) is essential for constructing a general DSD system.
@@
  • Not all network architectures are appropriate for the DSD task. As the good performance obtained from CNN-based network, ConvNeXt-Tiny, Whisper models, suggesting that CNN-based architectures are suitable for DSD task.
  • The ensemble of network architectures is effective in enhancing the model performance on the DSD task rather than the ensemble of spectrograms.
@@
  • Leveraging pre-trained models such as Whisper shows effectiveness, reinforcing the growing trend of using Encoder-Decoder architectures with pre-trained Encoders. This explains why these architectures have gained popularity in recent works.
In the cross-language evaluation, as shown in Table XII, all proposed DSD systems exhibit poor performance. This suggests that training a model on a single language (e.g., English) and testing it on other languages (e.g., French, German, Spanish, Italian) is not effective. To develop a robust DSD model for multiple languages, training with multilingual datasets is essential. This highlights the need for the DSD research community to focus on creating and publishing more multilingual datasets for the task.
TABLE XI: Performance comparison among deep learning models and ensemble of high-performance models
on ASVspoof 2021 (LA &DF Tasks) for cross-dataset evaluation
[…]
B5 STFT & LF Codec CNN ASV21-LA 0.84 0.87 0.89 0.16
B5 STFT & LF Mixup, Spec. CNN ASV21-LA 0.88 0.88 0.79 0.27
B5 STFT & LF Codec & Mixup, Spec. CNN ASV21-LA 0.85 0.87 0.90 0.15
B5 STFT & LF Codec CNN ASV21-DF 0.88 0.91 0.80 0.25
B5 STFT & LF Mixup, Spec. CNN ASV21-DF 0.91 0.88 0.77 0.28
B5 STFT & LF Codec & Mixup, Spec. CNN ASV21-DF 0.91 0.93 0.80 0.27
B3 CQT Mixup, Spec. CNN ASV21-LA 0.89 0.86 0.49 0.51
B5 STFT & LF Mixup, Spec. CNN ASV21-LA 0.88 0.88 0.79 0.27
B7 STFT & GAM Mixup, Spec. CNN ASV21-LA 0.89 0.87 0.52 0.49
B3 CQT Mixup, Spec. CNN ASV21-DF 0.95 0.94 0.51 0.49
B5 STFT & LF Mixup, Spec. CNN ASV21-DF 0.91 0.88 0.77 0.28
B7 STFT & GAM Mixup, Spec. CNN ASV21-DF 0.96 0.95 0.61 0.42
D7 STFT & LF Mixup, Spec. ConvNeXt-Tiny ASV21-LA 0.88 0.88 0.73 0.33
E1 Raw Audio None Whisper+MLP ASV21-LA 0.84 0.86 0.88 0.18
D7 STFT & LF Mixup, Spec. ConvNeXt-Tiny ASV21-DF 0.93 0.94 0.76 0.32
E1 Raw Audio None Whisper+MLP ASV21-DF 0.84 0.89 0.92 0.14
B3 + B5 + B7 CQT, STFT&LF, STFT&GAM Mixup, Spec. CNN ASV21-LA 0.90 0.87 0.75 0.30
D7 + E1 Raw Audio, STFT&LF Mixup, Spec. Whisper, CNN ASV21-LA 0.90 0.91 0.96 0.11
B3 + B5 + B7 CQT, STFT&LF, STFT&GAM Mixup, Spec. CNN ASV21-DF 0.96 0.95 0.77 0.29
D7 + E1 Raw Audio, STFT&LF Mixup, Spec. Whisper, CNN ASV21-DF 0.94 0.95 0.95 0.13
TABLE XII: Performance comparison among deep learning models and ensemble of high-performance models
on MLAAD dataset for cross-language evaluation
[…]
B5 STFT & LF Codec & Mixup, Spec. CNN MLAAD-DE 0.45 0.32 0.53 0.46
B5 STFT & LF Codec & Mixup, Spec. CNN MLAAD-IT 0.49 0.34 0.27 0.69
B5 STFT & LF Codec & Mixup, Spec. CNN MLAAD-FR 0.49 0.35 0.48 0.51
B5 STFT & LF Codec & Mixup, Spec. CNN MLAAD-ES 0.48 0.33 0.45 0.52
E1 Raw Audio None Whisper+MLP MLAAD-DE 0.53 0.52 0.56 0.45
E1 Raw Audio None Whisper+MLP MLAAD-IT 0.52 0.52 0.54 0.48
E1 Raw Audio None Whisper+MLP MLAAD-FR 0.59 0.57 0.62 0.40
E1 Raw Audio None Whisper+MLP MLAAD-ES 0.52 0.52 0.53 0.48
B5 + E1 Raw Audio, STFT & LF Codec & Mixup, Spec. CNN, Whisper+MLP MLAAD-DE 0.50 0.38 0.54 0.47
B5 + E1 Raw Audio, STFT & LF Codec & Mixup, Spec. CNN, Whisper+MLP MLAAD-IT 0.52 0.38 0.63 0.40
B5 + E1 Raw Audio, STFT & LF Codec & Mixup, Spec. CNN, Whisper+MLP MLAAD-FR 0.50 0.36 0.59 0.42
B5 + E1 Raw Audio, STFT & LF Codec & Mixup, Spec. CNN, Whisper+MLP MLAAD-ES 0.50 0.37 0.49 0.50